A charge of -3.96C is fixed at the center of a compass. Two additional charges are fixed on the circle of the compass (radius = 0.0946 m). The charges on the circle are -5.22C at the position due north and +6.27C at the position due east. What is (a) the magnitude and (b) direction of the net electrostatic force acting on the charge at the center? Specify the direction as an angle relative to due east.

east -W Fsouth
Answer #1

F1 = 9*10^9*3.96*5.22 / (0.0946)^2 (south)

F1 = 2.078*10^13 N

F2 = 9*10^9*3.96*6.27 / (0.0946)^2 (east)

F2 = 2.497*10^13 N

R = \small \sqrt{}(F1^2 + F2^2)

R = \small \sqrt{}((2.078*10^13)^2+(2.497*10^13)^2)

R = 3.248*10^13 N

θ = arctan (F1/F2)

θ = 39.76oC